OpenStudy (nbarrera):

Please Help Will Medal! Find the least common denominator of the two rational expressions:

OpenStudy (nbarrera):

Find the least common denominator of the two rational expressions: \[\frac{ 4k }{ k^{2} -9},\frac{ 2 }{ 3+k }\] A. The LCD is 3 – k. B. The LCD is k + 3. C. The LCD is k – 3. D. The LCD is k^2 – 9.

OpenStudy (nbarrera):

@welshfella

OpenStudy (nbarrera):

So can u help me @Brandonsnider222 ?

OpenStudy (anonymous):

3k is the correct answer 4/ k -9 -2/3 only terms in the denominator are k and 3 . LCD is their product 3k

OpenStudy (anonymous):

k - 9 = (k + 3)(k - 3) k + 3 = (k + 3) = (k + 3)(k - 3) = k - 3k + 3k - 9 = k - 9 Answer: (k + 3)(k - 3) OR k - 9 is the least common denominator.

OpenStudy (anonymous):

So B is the correct answer.

OpenStudy (nbarrera):

K Thx :)

OpenStudy (anonymous):

No problem